OpenStudy (anonymous):

The equation y = - 1777x + 27, 153 can be used to predict the number y of gun deaths in the United States x years after 2000, that is x = 0 corresponds to 2000, x = 3 corresponds to 2003, x = 5 corresponds to 2005, and so on. Predict the number of gun deaths in 2005 and 2007. In what year will number of gun deaths be 11,160? The predicted gun deaths in 2005 will be What? The predicted gun deaths in 2007 will be What? The predicted number of gun deaths will be 11,160 in the year?

OpenStudy (anonymous):

\[y(x)=y(2000+x)=27,153-1777x\] 1. Solve \[2000+x=2005\] For \(x\), then put \(x\) into \(27,153-1777x\). 2. Solve \[2000+x=2007\] For \(x\), then put \(x\) into \(27,153-1777x\). 3. Let \[y(2000+x)=y(x)=11,160\] Solve \[27,153-1777x=11,160\] For \(x\), then put \(x\) into \(2000+x\).